### v1

# Enhancing Test Time Adaptation with Few-shot Guidance

Abstract
Deep neural networks often encounter significant performance drops while facing with domain shifts between training (source) and test (target) data. To address this issue, Test Time Adaptation (TTA) methods have been proposed to adapt pre-trained source model to handle out-of-distribution streaming target data. Although these methods offer some relief, they lack a reliable mechanism for domain shift correction, which can often be erratic in real-world applications. In response, we develop Few-Shot Test Time Adaptation (FS-TTA), a novel and practical setting that utilizes a few-shot support set on top of TTA. Adhering to the principle of few inputs, big gains, FS-TTA reduces blind exploration in unseen target domains. Furthermore, we propose a two-stage framework to tackle FS-TTA, including (i) fine-tuning the pre-trained source model with few-shot support set, along with using feature diversity augmentation module to avoid overfitting, (ii) implementing test time adaptation based on prototype memory bank guidance to produce high quality pseudo-label for model adaptation. Through extensive experiments on three cross-domain classification benchmarks, we demonstrate the superior performance and reliability of our FS-TTA and framework.
Introduction
In recent years, deep neural networks have exhibited remarkable capabilities in representation learning. However, their performance relies heavily on the assumption that the distributions of training (source) and test (target) data are identical (Long et al. 2015; Ganin and Lempitsky 2015; Li et al. 2017). In real-world deployment, such a distribution shift is inevitable, as it is practically impossible to collect and annotate data for all possible environments in advance of training. Besides, this distribution shift can significantly degrade the performance of the deployed source model.
To address the aforementioned issues, numerous studies have proposed solutions via domain adaptation (Long et al. 2015; Tzeng et al. 2017; Zhang et al. 2019; Xiao and Zhang 2021; Xin et al. 2023) and domain generalization (Volpi et al. 2018; Zhou et al. 2021; Kim et al. 2021; Sicilia, Zhao, and Hwang 2023). While these approaches have demonstrated impressive performance gains on realistic benchmarks, a considerable gap remains between their problem settings and practical application scenarios. Domain adaptation relies on the impractical assumption that target domain data are available and participate in the source training process. In contrast, domain generalization aims to directly enhance the generalization of the source model without exploring the target domain data, even if they can be obtained during the test time.
Figure 1: Test Time Adaptation (TTA) vs. Few-Shot Test Time Adaptation (FS-TTA). FS-TTA utilizes labeled few-shot target data in addition to TTA. The results for TTA are derived from the performance of TENT (Wang et al. 2021) on the OfficeHome (Venkateswara et al. 2017).
In order to overcome these limitations of domain adaptation/generalization and protect the privacy of the source data, TENT (Wang et al. 2021) introduces fully test time adaptation (TTA). TTA aims to adapt a pre-trained source model to the target domain using input mini-batch data during the test time, without relying on source data or supervision. TTA is particularly focused on an online setting, where the model must adapt and make predictions immediately upon receiving each batch of potentially non-independent and identically distributed (non-i.i.d.) target samples. To serve this purpose, TENT employs test-time entropy minimization to reduce the generalization error on shifted target data. Additionally, extensive research has sought to improve TTA through various approaches such as pseudo-labeling (Iwasawa and Matsuo 2021; Wang, Zhang et al. 2023), consistency regularization (Boudiaf et al. 2022), and anti-forgetting regularization (Niu et al. 2022). While these methods can perform model adaptation during the test time, they encounter three primary challenges:
1) Domain shift correction: The certainty of TTA methods in addressing domain shifts effectively without utilizing target labels is questionable. The t-SNE visualization in Fig. 5 clearly illustrates this point, where we observe that the feature distribution exhibits negligible change following the adaptation process with TENT. This suggests that TTA methods may struggle to effectively adjust to new domain characteristics in the complete absence of target labels, which could provide essential guidance for adaptation.
2) Generalizability: The effectiveness of TTA methods varies across different scenarios. In some cases, they might even underperform compared to the pre-trained source model without any adaptation, as illustrated in Fig. 1 (Source Model vs. TENT). This variability indicates that the generalization performance of TTA methods is not particularly strong and can be influenced by various factors, including the domain shift and the specific characteristics of the model and dataset involved.
3) Data reliance: The success of TTA methods heavily relies on the availability and quality of unlabeled mini-batch data from the target domain. This reliance presents a challenge, as the adaptation process is directly influenced by the representativeness, quantity, and quality of the available unlabeled data. In scenarios where high-quality, relevant unlabeled data is scarce or not fully representative of the entire target domain, TTA methods may face difficulties in achieving optimal performance, highlighting a major limitation in their application across various real-world settings.
Figure 2: The performance impact of one-shot.
The fundamental reason for these challenges is the blind exploration of the target domain without any supervision information when domain shift exists. Actually, in practical scenarios, it’s quite feasible to obtain few-shot labeled samples from the target domain. Consequently, this leads us to ask: If given little target domain supervisory information, could the adaptation performance be improved?
To answer this question, we test the one-shot situation, as shown in Fig. 2. Specifically, we use one sample per class to fine-tune the source model with cross-entropy loss. We find that the performance is easily improved compared to TENT, which shows that little supervision information can be more effective than a large amount of unsupervised information.
Based on the aforementioned findings, we introduce Few-Shot Test Time Adaptation (FS-TTA), encapsulating the concept of few inputs, big gains. As illustrated in Fig. 1, by integrating a few-shot support set from the target domain prior to model adaptation, FS-TTA effectively reduces domain shift while retaining the source-free and online characteristics inherent to TTA. It is noted that such setting is particularly beneficial in scenarios where precision and reliability are paramount, such as medical image analysis and autonomous driving, even spending a few extra labeling costs.
To solve the FS-TTA, we develop an effective framework. For domain shift correction, we first fine-tune the pre-trained source model with the few-shot support set, fostering initial adaptation to the target. To prevent overfitting, we propose Feature Diversity Augmentation (FDA) to generate new features. During the test time, we employ a self-training strategy, which involves assigning pseudo-labels to unlabeled online mini-batches and using these labels to further update the model online. Furthermore, in order to reduce the impact of noisy pseudo-labels on the model, we propose Entropy Filter and Consistency Filter. The former filters out high-entropy samples with low confidence, and the latter is achieved through dual-branch prediction consistency. The experimental results across various cross-domain image recognition datasets show that our FS-TTA method significantly surpasses the performance of state-of-the-art methods and other baselines. To sum up, our main contributions are as follows:
  • New research direction: We highlight the setting of Few-Shot Test Time Adaptation (FS-TTA), in which an additional few-shot support set is available prior to test time adaptation. Thanks to the few-shot samples, domain shift can be reduced more effectively.
  • Novel framework: We propose a carefully crafted framework to tackle FS-TTA, including fine-tuning the pre-trained source model with feature diversity augmentation module and performing test time adaptation via high-quality pseudo-labeled samples in self-training manner.
  • Superior performance: We provide empirical results on various cross-domain classification benchmarks to demonstrate the superiority of our framework. In comparison to the state-of-the-art method in TTA, our method achieves improvements of 2.0% on PACS, 7.8% on OfficeHome, and 3.9% on DomainNet.
Related Work
Test Time Adaptation
Test Time Adaptation (TTA) strives to adapt the pre-trained source model during the test time to alleviate distribution shifts. Previous TTA methods address distribution alignment by incorporating self-supervised tasks (Sun et al. 2020). Although effective, TTA necessitates access to training samples and modifies the training process. To surmount this challenge, TENT (Wang et al. 2021) introduces the concept of fully test-time adaptation, utilizing only target data. TENT adapts the batch normalization layer by minimizing the entropy of model predictions. Distinct from TENT, recent works (Schneider et al. 2020; Nado et al. 2020) estimate the mean and variance of the activation in each batch normalization on the forthcoming test mini-batch. Consequently, LAME (Boudiaf et al. 2022) and EATA (Niu et al. 2022) concentrate on tackling the issue of catastrophic forgetting as the source model undergoes continuous updates. TSD (Wang, Zhang et al. 2023) integrates self-training into TTA, selecting high-quality test samples to fit the model.
Table 1: Comparison with various adaptation settings, where $s$ and $t$ denote source domain and target domain, respectively. $L^{d}$ and $U^{d}$ denote labeled datasets and unlabeled datasets from domain $d$. “Online” means that adaptation can predict a batch of incoming test samples immediately. “$k$” represents the number of samples per class. “C” indicates the number of classes for the target domain.
Setting Source-free Training inputs Online
Source domain(s) Target domain Size of available target data
[…]
Comparisons with Other Settings
We compare Few-Shot Test Time Adaptation (FS-TTA) with similar problem settings (details are in the appendix), as illustrated in Tab. 1.
  • Compared with Domain Generalization, FS-TTA eliminates the need for access to source data, thereby safeguarding the privacy of the source data. Moreover, FS-TTA facilitates adaptation to the downstream target domain via model parameter updates.
  • Compared with Source-Free Domain Adaptation, FS-TTA obviates the need to acquire all target domain data simultaneously and can dynamically update the model online, contingent upon incoming mini-batches.
  • Compared with Few-Shot Transfer Learning, FS-TTA can not only utilize a small number of target samples, but can also further update the model using online mini-batch target data during the test time.
  • Compared with Test Time Adaptation, FS-TTA capitalizes on an auxiliary small set of samples from the target, allowing the pre-trained source model to adapt more swiftly and proficiently to the target domain. Furthermore, FS-TTA demonstrates exceptional performance in managing circumstances involving substantial domain shifts.
Method
Problem Setting
Considering a typical scenario where a source model $f_{\theta_{s}}$, is equipped with parameters $\theta_{s}$ and trained on source datasets $\mathcal{D}_{s_{1}},\mathcal{D}_{s_{2}},...,\mathcal{D}_{s_{n}}$, our goal is to adapt this pre-trained source model to a target domain $D_{t}$ without accessing source data. We postulate the existence of a small, labeled support set $S=\{(s_{i},y_{i})\}$ derived from $D_{t}$, where $s_{i}$ represents the image, and $y_{i}$ is its corresponding label. During the test time, mini-batches unlabeled target samples online arrive. Few-Shot Test Time Adaptation (FS-TTA) strives to effectively adapt the pre-trained source model $f_{\theta_{s}}$, utilizing the labeled support set $S$ and unlabeled online mini-batch to tackle the domain shift challenge. Notably, the support set $S$ could be offline obtained before the test time.
Stage I: Model Adaptation via Fine-Tuning
To significantly and swiftly enhance the initialization performance of the pre-trained source model in the target domain and minimize domain shifts, we design to fine-tune the pre-trained source model using the few-shot support set. Given the limited number of samples per class, there is a potential risk of overfitting during the fine-tuning process. To mitigate this, we introduce the Feature Diversity Augmentation (FDA) module, which generates new features by mixing statistics. Ultimately, we use a supervised classification loss to fine-tune the pre-trained source model. This entire procedure is illustrated in Stage I of Fig. 3.
Figure 3: Illustration of our two-stage framework. In Stage I, we employ the few-shot support set to fine-tune the source model. To prevent overfitting, we propose FDA module. In Stage II, we maintain a prototype memory bank to guide test time adaptation. In order to update the prototype memory bank and model with effective samples, we propose the entropy filter and consistency selection modules.
Feature Diversity Augmentation.
Prior research (Zhou et al. 2021) has demonstrated a significant association between feature statistics and image style, which is intricately linked to data distribution within the field of computer vision. Inspired by their discoveries, we propose Feature Diversity Augmentation (FDA), which enhances feature diversity and mitigates the risk of overfitting during fine-tuning.
FDA is incorporated between layers (blocks) in the pre-trained source backbone, as depicted in Fig. 3. More specifically, FDA mixes the feature statistics of two random samples to generate new features. The computations within FDA module can be summarized in three steps. Firstly, given two feature maps $f_{i}$ and $f_{j}$ from the support set, we compute their feature statistics ($\mu_{i},\sigma_{i}$) and ($\mu_{j},\sigma_{j}$). Secondly, FDA generates the mixtures of feature statistics:
$\boldsymbol{\gamma}_{\text{mix }}=\lambda\sigma_{i}+(1-\lambda)\sigma_{j},$ (1)
$\boldsymbol{\beta}_{\text{mix }}=\lambda\mu_{i}+(1-\lambda)\mu_{j}.$ (2)
In this case, $\lambda$ denotes the mixing ratio coefficient. Ultimately, the mixtures of feature statistics are applied to the feature map $f_{i}$ via instance normalization:
$f_{i}^{\prime}=\boldsymbol{\gamma}_{\operatorname{mix}}\odot\frac{f_{i}-\mu_{i}}{\sigma_{i}}+\boldsymbol{\beta}_{\operatorname{mix}},$ (3)
where $f_{i}^{\prime}$ represents the newly generated feature map.
Fine-Tuning Source Model.
To enhance the adaptation of the pre-trained source model to the target, we employ the few-shot support set to fine-tune the model with the FDA module. Specifically, the few-shot support set is processed through $f_{\theta_{s}}$ to minimize a supervised loss, defined as:
$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{k*C}\mathcal{H}\left(y_{i},p\left(\hat{y_{i}}\mid\boldsymbol{s}_{i}\right)\right),$ (4)
where $\mathcal{H}(\cdot)$ is the cross-entropy loss. The term $y_{i}$ is the ground-truth label of $s_{i}$, indicating one of sample from few-shot support set, and $C$ represents categories of the target.
Table 2: Compared with test time adaptation methods on three datasets with ResNet50 backbone.
OfficeHome PACS DomainNet
Method Art Clip Prod Real Avg. Art Cart Phot Sket Avg. Avg.
Test time adaptation methods
ERM (Vapnik 1999) 60.7 55.7 76.2 76.8 67.4 82.5 80.8 94.0 80.9 84.5 45.2
BN (Nado et al. 2020) 58.2 55.6 75.1 75.5 66.1 83.2 84.9 94.0 77.9 85.0 43.3
TENT (Wang et al. 2021) 60.6 58.7 76.5 76.1 68.0 85.2 86.7 94.9 82.9 87.4 44.7
T3A (Iwasawa and Matsuo 2021) 61.2 56.7 78.0 77.3 68.3 84.0 82.3 95.0 82.7 86.0 46.1
ETA (Niu et al. 2022) 58.4 55.8 75.2 75.5 66.2 83.2 84.9 94.0 77.9 85.0 46.1
LAME (Boudiaf et al. 2022) 58.7 55.6 75.1 75.4 66.2 84.9 85.5 95.0 80.9 86.6 43.2
TSD (Wang, Zhang et al. 2023) 62.3 57.5 77.5 77.5 68.7 87.6 88.7 96.1 85.0 89.4 47.7
Fine-tuning + Test time adaptation methods
FT+TENT (Wang et al. 2021) 68.8 65.5 79.8 78.5 73.2 87.0 86.9 95.2 83.6 88.2 45.4
FT+TSD (Wang, Zhang et al. 2023) 70.5 65.1 80.3 79.2 73.8 88.3 88.6 96.5 85.9 89.8 48.5
FS-TTA 73.2 68.3 83.0 81.6 76.5 90.4 89.7 97.6 87.8 91.4 51.6
$\Delta_{up}$ over TSD (+10.9$\uparrow$ (+10.8$\uparrow$ (+5.5$\uparrow$ (+4.1$\uparrow$ (+7.8$\uparrow$ (+2.8$\uparrow$ (+1.0$\uparrow$ (+1.5$\uparrow$ (+2.8$\uparrow$ (+2.0$\uparrow$ (+3.9$\uparrow$
Stage II: Test Time Adaptation
During this stage, a mini-batch of unlabeled samples, denoted as $x=\{x_{1},x_{2},..,x_{B}\}$, online arrives. The central concept of this stage is to employ a self-training strategy to update the fine-tuned source model online, enabling it to fully adapt to the target domain. This involves assigning pseudo-labels to unlabeled online mini-batches and using these labels to further update the model. Thus, we first generate the pseudo-labels by $\hat{y_{i}}=\text{argmax}(p_{i})$ for $x_{i}$, where $p_{i}$ is the prediction logits. However, it is inevitable that there are always some noisy samples are misclassified, leading to wrong pseudo-labels. To address this issue, we propose two modules to produce high quality pseudo-labels. The first is entropy filter, which screens out unreliable samples using Shannon entropy (Shannon 2001). Typically, samples with higher entropy are considered to have lower prediction confidence. The second module is a prototype memory bank classification, which works in tandem with the classifier. The prototype memory bank is used to generate pseudo-labels outside the classifier, according to the nearest class prototype in the feature space. After that, pseudo-labels with consistency prediction is preserved for model adaptation. The entire process is outlined in Stage II of Fig. 3.
Entropy Filter.
To dynamically update the model using online mini-batch target, it is crucial to filter out noisy samples, as they may be assigned to incorrect classes, resulting in inaccurate prototype computation. In this regard, we propose the Entropy Filter, which employs Shannon entropy (Shannon 2001) to select confident samples in the mini-batch. For an sample $x_{i}$, its entropy can be computed as:
$H(p_{i})=-\sum(p_{i})\cdot log(p_{i}).$ (5)
Based on the insights from previous work (Wang et al. 2021), high entropy samples should be filtered out, as lower entropy typically indicates higher accuracy. Consequently, we sort the entropy of all samples in the mini-batch and select the top $\alpha\%$ samples with lower entropy, donated as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},...,\hat{x}_{|\alpha\%*B|}\}$.
Prototype Memory Bank.
We maintain a prototype memory bank $M=\{m_{1},m_{2},...,m_{C}\}$ to store class prototypes, where $C$ represents categories of the target. The prototype memory bank is initialized with the few-shot support set $S$, defined as:
$m_{c_{0}}=\frac{\sum_{i=1}^{|S|}f_{i}\cdot\mathbbm{1}[y_{i}=c]}{\sum_{i=1}^{|S|}\mathbbm{1}[y_{i}=c]},$ (6)
where $\mathbbm{1}[\cdot]$represents an indicator function, yielding a value of 1 if the argument is true or 0 otherwise, and $m_{c_{0}}$ denotes the initial moment of the $c$-th class prototype. Thanks to the few-shot support, precise guidance can be provided during the initial phase, thereby reducing reliance on the quality of online mini-batch data.
Throughout the test time adaptation process, we persistently update the prototype memory bank by incorporating selected reliable samples with pseudo labels:
$m_{c_{t}}=\beta*m_{c_{t-1}}+(1-\beta)*\frac{\sum_{j=1}^{|\hat{x}|}f_{j}\cdot\mathbbm{1}[\hat{y}_{j}=c]}{\sum_{j=1}^{|\hat{x}|}\mathbbm{1}[\hat{y}_{j}=c]},$ (7)
where $m_{c_{t}}$ represents the $c$-th class prototype at time $t$, and $\beta$ represents the sliding update coefficient.
Test Time Adaptation.
During the test time adaptation, we adopt high-quality pseudo-labeled samples to guide the model update. First, we define the prototype-based classification output as the softmax over the feature similarity to prototypes for class $c$:
$\hat{p}_{j}^{c}=\frac{\exp\left(\operatorname{sim}\left(f_{j},m_{c}\right)\right)}{\sum_{c=1}^{C}\exp\left(\operatorname{sim}\left(f_{j},m_{c}\right)\right)},$ (8)
where $sim(\cdot,\cdot)$ represents cosine similarity. Subsequently, we propose that, for a reliable sample, the outputs of the fine-tuned model and prototype-based classification should be similar. Therefore, we propose the consistency filter to identify incorrect predictions. This strategy can be implemented through a filter mask for samples $x_{j}$ as follows:
$\mathcal{M}_{j}=\mathbbm{1}\left[\arg\max p_{j}=\arg\max\hat{p}_{j}\right].$ (9)
Ultimately, we can update the model using reliable samples, and the loss can be formulated as follows:
$\mathcal{L}_{online}=\frac{{\textstyle\sum_{j=1}^{|\hat{x}|}}\mathcal{H}_{j}*\mathcal{M}_{j}}{{\textstyle\sum_{j=1}^{|\hat{x}|}}\mathcal{M}_{j}}.$ (10)
It’s noteworthy that our self-training process does not involve specifying any threshold, which enhances the model’s generalizability.
Experiment
Experimental Settings
Dataset.
To test the effectiveness of our setting and method, we experiment on three cross-domain benchmarks. PACS (Li et al. 2017) consists of four distinct domains: Art, Cartoon, Photo, and Sketch, and comprises a total of 9,991 images and 7 classes. Office-Home (Venkateswara et al. 2017) consists of 15,588 images from four distinct image domains, namely Real World, Clipart, Art, and Product, with 65 classes. DomainNet (Peng et al. 2019) is a large-scale dataset that comprises six domains, denoted as $d\in${Clipart, Infograph, Painting, Quickdraw, Real, Sketch}, with a total of 586,575 images and 345 classes.
Implementation Details.
In the main experiments, we utilize ResNet-50 (He et al. 2016) pre-trained on ImageNet-1k (Russakovsky et al. 2015) as our backbone, which is widely used in test time adaptation literature. For source model training, we adhere to the leave-one-domain-out protocol, as recommended by previous work  (Wang, Zhang et al. 2023; Zhou et al. 2021). We employ the Adam optimizer with a learning rate of $5e^{-5}$. For few-shot test time adaptation, we also employ the Adam optimizer and set the batch size to The few-shot support set typically selects 5 to 16 samples per class, depending on the difficulty of the target. We carry out all experiments on NVIDIA V100 GPUs.
Baselines.
We compare our method with test time adaptation methods, BN (Nado et al. 2020), TENT (Wang et al. 2021), ETA (Niu et al. 2022), T3A (Iwasawa and Matsuo 2021), LAME (Boudiaf et al. 2022), and TSD (Wang, Zhang et al. 2023). Additionally, we create new baselines by combining fine-tuning with TTA methods for a more comprehensive comparison. We also compare with some methods in domain generalization and source-free domain adaptation, including DNA (Chu et al. 2022), PCL (Yao et al. 2022), SWAD (Cha et al. 2021), and F-mix (Kundu et al. 2022).
Table 3: Compared with existing DG and SFDA methods on OfficeHome and DomainNet.
OfficeHome DomainNet
Method Art Clip Prod Real Avg. Clip Info Pain Quic Real Sket Avg.
Domain generalization methods
ERM (Vapnik 1999) 60.7 55.7 76.2 76.8 67.4 64.8 22.1 51.8 13.8 64.7 54.0 45.2
DNA (Chu et al. 2022) 67.7 57.7 78.9 80.5 71.2 66.1 23.0 54.6 16.7 65.8 56.8 47.2
PCL (Yao et al. 2022) 67.3 59.9 78.7 80.7 71.6 67.9 24.3 55.3 15.7 66.6 56.4 47.7
SWAD (Cha et al. 2021) 66.1 57.7 78.4 80.2 70.6 66.1 22.4 53.6 16.3 65.5 56.2 46.7
Source-free domain adaptation methods
F-mix (Kundu et al. 2022) 72.6 67.4 85.9 83.6 77.4 75.4 24.6 57.8 23.6 65.8 58.5 51.0
FS-TTA 73.2 68.3 83.0 81.6 76.5 68.6 30.8 56.4 24.2 69.1 60.2 51.6
SWAD + FS-TTA 77.4 71.1 86.4 84.2 79.8 - - - - - - -
Performance Comparisons
Comparison with TTA methods.
Tab. 2 details the comparison results between our method and various TTA methods on the Office-Home and PACS datasets, as well as the final results of DomainNet (detailed in Tab. 3). We observe that our method achieves state-of-the-art performance.
Primarily, our approach exhibits a significant enhancement in performance compared to the source model (ERM). Our FS-TTA achieves improvements across all four tasks on Office-Home, with gains of 12.5% (Art), 12.6% (Clipart), 6.8% (Product), and 4.8% (Real), respectively. Notably, our method demonstrates more substantial improvement on the more challenging tasks (e.g., Art and Clipart), confirming that FS-TTA is more friendly for large domain shifts. On the other two datasets, we observe average performance increments of 6.9% (PACS) and 6.4% (DomainNet).
Moreover, our method outperforms the state-of-the-art TTA method, TSD, with average performance increments of 2.0% (PACS), 7.8% (Office-Home), and 3.9% (DomainNet). The lesser improvement in PACS can be attributed to its lower complexity, while our method shows superior performance on the more challenging Office-Home and DomainNet datasets. This significant improvement benefits from our effective utilization of few-shot target information, including the FDA module and initializing the prototype memory bank. The performance of some TTA methods, such as ETA and LAME, does not meet the expected standards on Office-Home and other datasets. In fact, they even exhibit inferior performance compared to the source model on certain tasks (e.g., Art, Product, and Real), which highlights the limitations of TTA and the necessity of few-shot target samples. In conclusion, our FS-TTA demonstrates a notable advantage in tasks that closely resemble real-world scenarios and provides a significant boost in performance with minimal additional computational overhead.
Finally, for a more comprehensive comparison with TTA methods, we construct new baselines, which combine fine-tuning with TTA methods (e.g., TENT and TSD). According to the results in Tab. 2, our method shows an average (three datasets) improvement of 4.2% over Fine-Tuning+TENT, and 2.4% over Fine-Tuning+TSD, which demonstrates the superiority of our method over existing TTA techniques in migrating to few-shot TTA scenarios. This result benefits from our unique FDA module design in Fine-Tuning and the use of few-shot support set to initialize prototype memory bank.
Refer to captionRefer to caption(a) Different stage(b) Different $\alpha$
Figure 4: (a) effectiveness analysis about two-stage framework and (b) sensitivity analysis about parameter $\alpha$.
Comparison with DG/SFDA methods.
The above experiments mainly focus on TTA, which aims to adapt the model during the test time. A natural question arises: How about our method compared with domain generalization (DG) or source-free domain adaptation (SFDA) methods?
To answer this question, we compare our method with several methods in DG and SFDA. The results of Office-Home dataset are shown in Tab. 3. It can be seen that our method outperforms the state-of-the-art methods in DG, such as SWAD and PCL. Furthermore, equipped with SWAD (SWAD+FS-TTA), our method achieves 79.8% accuracy. This result benefits from our adaptation of the model during the test time. In comparison to advanced SFDA methods, FS-TTA still achieves satisfactory results. It is worth noting that FS-TTA is more flexible in real-world scenarios than SFDA since it adapts the target data in an offline manner, requiring more training loops and resources. The results of DomainNet are shown in Tab. 3. The overall performance of FS-TTA outperforms the SFDA methods, suggesting that FS-TTA is more adept at handling challenging tasks.
Ablation Study
Figure 5: The t-SNE feature visualization of (a) ERM, (b) TENT, (c) LAME, and (d) our method.
Effectiveness of two-stage framework.
Our proposed method consists of two stages, with the individual contributions of each stage presented in Fig. 4(a). Compared to the baseline source model, Stage I of our approach achieves an average improvement of 6.6% on the Office-Home. This highlights the effectiveness of our fine-tuning strategy, which employs a mixture of statistics between samples, validating its suitability for the target domain. Our test time adaptation method, which relies on class prototype memory bank guidance during Stage II, adds an extra 2.5% performance enhancement. As a result, our two-stage framework establishes itself as a robust foundation for the Few-Shot Test Time Adaptation setting, demonstrating its considerable potential in enabling online model adaptation in real-world situations where labeled data is scarce.
Sensitivity to $\alpha$.
The parameter $\alpha$ represents the proportion of each batch that is selected through an entropy filter to update the prototype memory bank and the model. To evaluate the impact of $\alpha$, we conduct an experimental analysis on the Office-Home dataset by assigning $\alpha$ to 0, 0.3, 0.6, and 1, respectively. The results, as shown in Fig. 4(b), demonstrate that $\alpha>0$ yields performance improvements compared to $\alpha=0$ (the source model), highlighting the effectiveness of our proposed framework. Furthermore, $\alpha=0.3$ and $\alpha=0.6$ perform better than $\alpha=1$ (no filter), indicating the effectiveness of our entropy filter strategy.
Qualitative analysis by t-SNE visualization.
We present t-SNE visualizations to compare the feature representations of the pre-trained source model (ERM), test time adaptation methods (TENT and LAME), and our proposed method, as illustrated in Fig. 5. The learned features of the pre-trained source model on the target domain are not well-separated due to the significant domain gap, as shown in Fig. 5(a). Additionally, we can observe no considerable feature distribution changes on the target domain after adaptation with TENT and LAME methods, as shown in Fig. 5(b) and Fig. 5(c). In contrast, our method produces more uniform and aligned feature distribution after adapting to the target domain, as shown in Fig. 5(d).
Efficiency analysis.
In our main experiments, we opt for a mini-batch size of 64. To examine the variations in performance and computational efficiency with different batch size during test-time adaptation, we conduct a series of analytical experiments. As shown in Fig. 6(a), we observe that accuracy experiences a gradual increase as the batch size incrementally grows, reaching a plateau around a batch size of 64. In contrast, as shown in Fig. 6(b), running time exhibits a decreasing trend as the batch size grows. However, beyond a batch size of 64, the running time appears to stabilize. Consequently, for real-world applications aiming to achieve a trade-off between accuracy and computational efficiency, we suggest a batch size in the vicinity of 64.
Refer to captionRefer to caption(a) Accuracy (batch size)(b) Running Time
Figure 6: During the test time adaptation, we analyze:(a) accuracy and (b) running time, across various batchsize.
Figure 7: Ablation experiments on shot size.
Ablation experiments on shot size.
To elucidate the impact of the number of k-shots on our method, we carry out additional ablation experiments within the Office-Home dataset. The findings, illustrated in Figure 7, indicate a significant performance enhancement when the shot size ranges from 1 to 10, demonstrating a rapid performance ascension in this few-shot regime. Remarkably, even minimal shot sizes such as 1-shot and 3-shot exhibit substantial effectiveness. For instance, the 3-shot configuration achieves a 3.8% performance improvement over the TSD.
Conclusion
In this work, we introduce Few-Shot Test Time Adaptation (FS-TTA), a novel setting that diverges from traditional TTA by leveraging the few-shot support set to improve adaptation to the target. To tackle FS-TTA, we propose an effective framework, which involves employing the few-shot support set to fine-tune the pre-trained source model and maintaining a prototype memory bank to guide the test time adaptation. Results on three cross-domain benchmarks demonstrate the superior performance and reliability of our method. Looking ahead, we aspire to expand FS-TTA beyond current scope by investigating potential real-world tasks, instead of limiting to image recognition.